OpenStudy (anonymous):

3n+9+4n=12 (Show Work)

OpenStudy (karatechopper):

add the like terms together.. 3n+9+4n=12 3n+4n+9=12 7n+9=12 7n+9-9=12-9 7n=3 7n/7=3/7 n=3/7